OpenStudy (anonymous):

PLEASE HELP!!!! The photic zone describes the region in which microscopic plant life receive enough light to survive. The photic zone goes to a depth where about 1% of the surface light still remains. Find the depth of the photic zone if the intensity of light (L) at a depth (d) in meters is given by L=L0e^(-0.028d), where L0 is the intensity of surface light. You may use the graphing part of your calculator for this question. Possible answers are: a. 124 m b. 134 m c. 144 m d. 154 m e. 164 m

OpenStudy (anonymous):

I know you can use your calculator to find it but I have no idea how!!!

OpenStudy (anonymous):

L=100 per cent LO=1 per cent

OpenStudy (anonymous):

Ok, so how can I use that to solve? :( do I plug 0.01e^(-0.028x) into my calculator???

OpenStudy (anonymous):

You have an equation, the given, as they say. It has three unknowns -- not good. We have found two of the unknowns -- better. So now with one unknown you can solve for d.\[100percent=1percente ^{-0.028d}\]

OpenStudy (anonymous):

That makes much more sense, thank you. However im not so great at solving with exponents. How do I solve? This is how I would do it but I have noooo idea. 100 10,000 ---- = -0.028d ===> ------- = d ===> d= -357142.86 0.01 -0.028 I know its wrong but idk why :(

OpenStudy (anonymous):

You lost your e. You can't just arbitrarily take away your e. The plants would die from lack of sunlight.

OpenStudy (anonymous):

100 ---- = -0.028d should it start like that? 0.01e

OpenStudy (anonymous):

\[100=e ^{-0.028d}\]You have to use logarithms. Who knew marine biology was so much work.\[\ln 100=\ln e ^{-0.028d}\]

OpenStudy (anonymous):

So we just forget about the 0.01? does it turn into 100 ------ = d? that gives me d=-3571.43 -0.028

OpenStudy (anonymous):

There is no 0.01. The percentage signs cancel leaves 100 over one which is 100

OpenStudy (anonymous):

Ok, but I still solved and got -3571.43, which is not one of my answers. I dont understand!

OpenStudy (anonymous):

May be you did something wrong\[\ln 100=- 0.028d \ln e \]

OpenStudy (anonymous):

which gives me ln (100/-0.028)???? Im sorry I dont understand how to do these at all and I need to solve this problem

OpenStudy (anonymous):

You have the ability to find ln 100 and ln e on your calculator

OpenStudy (anonymous):

lne = 1 and ln100 =4.6, correct???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so then I take 4.6/-0.028= -164.28, correct?

OpenStudy (anonymous):

Yes but depth is always positive, so absolute value (erase negative sign)

OpenStudy (anonymous):

So the answer is E, 164 m? Thanks

OpenStudy (anonymous):

You're pretty good.